OpenStudy (anonymous):

How do I solve -3c-12=-5+c

OpenStudy (astrophysics):

Collect the like terms, so move all the things with c on one side and all the numbers on the other

OpenStudy (astrophysics):

Then you can solve for c alone, so see if you can set it up

OpenStudy (astrophysics):

-3x-12=-5+c so have can we move the +c on the right hand side to the left side?

OpenStudy (anonymous):

-2c=-17?

OpenStudy (astrophysics):

-3c-c=-5+12

OpenStudy (astrophysics):

What is -3c-c?

OpenStudy (anonymous):

-4

OpenStudy (astrophysics):

Good, so -4c=-5+12 now what is -5+12?

OpenStudy (anonymous):

7

OpenStudy (astrophysics):

Yup, so -4c=7 now we just have to divide both sides by -4 and you'll have c

OpenStudy (anonymous):

c=-1.75?

OpenStudy (astrophysics):

Yes, or c=-7/4